# Self-consistent constraints on the collision terms in quantum kinetic theory

### v1

@@ Abstract
We derive self-consistent constraint conditions for collision terms in quantum kinetic theory within Wigner function formalism. We present simple specific solutions for the collision terms in quantum kinetic theory which is compatible with these constraints. We obtain the self-consistent quantum kinetic theory at relaxation-time approximation, with which we discuss spin polarization effects and electric charge separation effects in relativistic heavy-ion collisions.
1 Introduction
@@
Collision terms in kinetic equation play a central role in describing the evolution from non-equilibrium to equilibrium, and can be derived from the BBGKY equation or hierarchyBogolyubov:1946 by introducing certain additional assumptions. As we all know, in general, the collision terms should satisfy the principle of detailed balance and the constraints of charge conservation or energy-momentum conservation. In this paper, we will demonstrate that the collision terms in quantum kinetic theory must satisfy some extra constraints totally determined by the self-consistent nature of the quantum kinetic equation. With this self-consistent constraints, we will give non-trivial specific solutions for the collision terms in quantum kinetic theory. We put the specific solution in the form of relaxation-time approximation and use it to discuss spin polarization and electric charge separation in relativistic heavy-ion collisions.
@@
2 Wigner function formalism
@@
3 Self-consistent constraints on collision terms
@@
3.1 Zeroth-order result
@@
3.2 First-order result
@@
At first order, we have obtained the constraints for the collision functions and the kinetic equation for zeroth-order independent Wigner functions ${\mathscr{J}}_{s,n}^{(0)}$ and ${\mathscr{M}}^{(0)}_{\mu}$. Since these constraint equations are derived only from the consistency of the Wigner equations, we call them self-consistent constraint equations.
3.3 Second-order result
@@
Following the same procedure as we have taken at first order, we can obtain the self-consistent constraints for the collision terms at the second order. Besides much more complicated vector algebraic and derivative operation at second order, we also encounter the terms associated with the derivative terms with $n^{\mu}$ such as $\partial_{x}^{\nu}n^{\mu}$ in the second-order constraint equations. However, it is remarkable that all these terms cancel each other if we impose the first-order constraint equations (91)-(95). The final independent constraint equations are given by
@@
4 Self-consistent solutions for collision terms
As we all know, these collision terms cannot be determined by a group of closed equations due to BBGKY hierarchy. Some proper approximation must be imposed to make the equations closed. No matter what approximation we make, the self-consistent constraints derived in previous sections should be fulfilled. In this section, we will find self-consistent specific solutions for these collision terms.
Let us start from the first-order constraints conditions (91)-(95) and try to find a specific expression as simple as possible but still not very trivial. In the simplest case, we will assume that all the collision terms at the first order are on-shell. It is easy to verify that the following expressions always satisfy all first-order constraints conditions (91)-(95),
@@ 4 Self-consistent solutions for collision terms
From the second-order constraints (122)-(126), we can find a second-order solution
@@
5 Self-consistent relaxation-time approximation
@@ 6 Summary and discussion
In this paper, we start from the gauge-invariant Wigner function formalism in quantum electrodynamics and derive the quantum kinetic theory with BBGKY hierarchy or general collisions terms. With the help of the semiclassical $\hbar$ expansion, we use the property of self-consistency of the Wigner equations to constrain the collision terms up to second order. These constraint equations provide general necessary conditions to make some approximation or simplification in some specific cases just like we impose the conservation laws in the collision terms. When only three vector functions involved, we present a specific solution for the collision terms in the formalism of GCKT. We further specialize this solution in the form of relaxation-time approximation and present the self-consistent quantum kinetic theory at relaxation-time approximation which is compatible with the self-consistent constraints. We find the quantum kinetic theory at relaxation-time approximation can be greatly simplified by defining or decomposing the distribution functions in the comoving fluid frame with velocity $u^{\mu}$. Some Wigner functions exhibit even no explicit dependence on the collision terms and the implicit dependence is totally coded into the distribution functions when we write the equations or functions in the fluid comoving frame.
@@
Although we didn’t take into account the mean electromagnetic field during deriving the quantum kinetic equation, Wigner functions, and the constraint conditions with collision terms, it is straightforward and easy to obtain the results when a mean field is involved. As we mentioned in the section 2, the collision functions discussed in this work actually contain all possible terms, including mean field contribution. At mean field approximation, all the constraint conditions derived in previous sections will be satisfied automatically as verified in Gao:2018wmr ; Gao:2019znl ; Ma:2022ins . Therefore, we can simply decompose the general electromagnetic field in (5) into mean field part and quantized field part, separate the mean field part from the quantized field, and directly obtain the quantum kinetic equation and Wigner functions with mean electromagnetic field involved. The constraint conditions for collision terms from the quantized field part remain unchanged.
@@
We have presented the quantum kinetic equations and Wigner equations in 8-dimensional phase space, 4-dimensional coordinate space $x^{\mu}$ plus 4-dimensional momentum space $p^{\mu}$. In 8-dimensional phase space, the on-shell Dirac delta function is always involved. We can integrate the time-like component of momentum to eliminate the singular Dirac delta function and obtain the quantum kinetic theory in 7-dimensional phase space, which can be applied to make numerical calculation directly. Actually, it is a trivial task to obtain the 7-dimensional quantum kinetic theory from the 8-dimensional one in sections 4 and 5 because only the onshell Dirac delta functions are involved and there are no derivative terms with respect to the momentum when the background electromagnetic field is absent.

### v2

@@ Abstract
We derive self-consistent constraint conditions for collision terms in quantum kinetic theory using the Wigner function formalism. We present specific solutions for these collision terms that align with the constraints. we develop quantum kinetic theory at relaxation-time approximation. With this framework, we discuss spin polarization and electric charge separation effects in relativistic heavy-ion collisions.
I Introduction
@@
Collision terms in kinetic equation play a central role in describing the evolution in non-equilibrium systume, and can be derived from the BBGKY equation or hierarchyBogolyubov:1946 by introducing certain additional assumptions. As we all know, in general, the collision terms should satisfy the principle of detailed balance and the constraints of charge conservation or energy-momentum conservation. In this paper, we will demonstrate that the collision terms in quantum kinetic theory must satisfy some extra constraints totally determined by the self-consistent nature of the quantum kinetic equation. With this self-consistent constraints, we will give non-trivial specific solutions for the collision terms in quantum kinetic theory. We put the specific solution in the form of relaxation-time approximation and use it to discuss spin polarization and electric charge separation in relativistic heavy-ion collisions.
@@
II Wigner function formalism
@@
III Self-consistent constraints on collision terms
@@
III.1 Zeroth-order result
@@
III.2 First-order result
@@
At first order, we have obtained the constraints for the zeroth-order collision functions and the kinetic equation for zeroth-order independent Wigner functions ${\mathscr{J}}_{s,n}^{(0)}$ and ${\mathscr{M}}^{(0)}_{\mu}$. Since these constraint equations are derived only from the consistency of the Wigner equations, we call them self-consistent constraint equations.
III.3 Second-order result
@@
Following the same procedure as we have taken at first order, we can obtain the self-consistent constraints for the collision terms at the first order. Besides much more complicated vector algebraic and derivative operation at second order, we also encounter the terms associated with the derivative terms with $n^{\mu}$ such as $\partial_{x}^{\nu}n^{\mu}$ in the second-order constraint equations. However, it is remarkable that all these terms cancel each other if we impose the zeroth-order constraint equations (91)-(95). The final independent constraint equations are given by
@@
IV Some other constraint conditions
In addition to the constraint equations for zeroth order (91)-(95) and first order (122)-(126) which are derived from the self-consistency of the Wigner equations, we can obtain other constraint equations from the specific structure defined from (22) to (29). The antisymmetry of the electromagnetic field tensor $F^{\mu\lambda}$ requires the following constraints
$\partial_{\nu}^{p}{\mathscr{C}^{\nu}}$ $=$ $0,\ \ \ \partial_{\nu}^{p}{\mathscr{C}^{\nu}_{5}}=0,\ \ \ \partial_{\mu}^{p}\mathscr{C}^{\mu\nu}_{s}=0,\ \ \ \partial_{\nu}^{p}{\mathscr{C}^{\nu\alpha\beta}}=0,$ (127)
$\partial_{\nu}^{p}\Delta{\mathscr{C}^{\nu}}$ $=$ $0\displaystyle 0,\ \ \partial_{\nu}^{p}\Delta{\mathscr{C}^{\nu}_{5}}=0,\ \ \partial_{\mu}^{p}\Delta\mathscr{C}^{\mu\nu}_{s}=0,\ \ \partial_{\nu}^{p}\Delta{\mathscr{C}^{\nu\alpha\beta}}=$ 0 . (128)
The constraints (127) require that the zeroth-order ${\mathscr{C}^{(0)\nu}}$, ${\mathscr{C}^{(0)\nu}_{5}}$, $\mathscr{C}^{(0)\mu\nu}_{s}$ and ${\mathscr{C}^{(0)\nu\alpha\beta}}$ in (91)-(95) should satisfy the same constraints. However, the constraints (127) and (128) cannot lead to the similar constraints on the first-order ${\tilde{\mathscr{C}}^{(1)\nu}}$, ${\tilde{\mathscr{C}}^{(1)\nu}_{5}}$, $\tilde{\mathscr{C}}^{(1)\mu\nu}_{s}$ and $\tilde{\mathscr{C}}^{(1)\nu\alpha\beta}$ in (127) and (128) due to the out-of-step linear combination in the definitions (108)-(111). In the following sections, we will disregard these constraints and only focus on the constraints (91)-(95) and (122)-(126).
V Self-consistent specific solutions for collision terms
As we all know, these collision terms cannot be determined by a group of closed equations due to BBGKY hierarchy. Some proper approximation must be imposed to make the equations closed. No matter what approximation we make, the self-consistent constraints (91)-(95) and (122)-(126) derived in previous sections should be fulfilled. In this section, we will find self-consistent particular solutions for these collision terms.
Let us start from the zeroth-order constraints conditions (91)-(95) and try to find a specific expression as simple as possible but still not very trivial. In the simplest case, we will assume that all the collision terms at the first order are on-shell. It is easy to verify that the following expressions always satisfy all first-order constraints conditions (91)-(95),
@@ V Self-consistent specific solutions for collision terms
From the second-order constraints (122)-(126), we can find a first-order solution
@@
VI Self-consistent relaxation-time approximation
@@ VII Summary and discussion
In this paper, we start from the gauge-invariant Wigner function formalism in quantum electrodynamics and derive the quantum kinetic theory with BBGKY hierarchy or general collisions terms. With the help of the semiclassical $\hbar$ expansion, we use the property of self-consistency of the Wigner equations to constrain the collision terms up to second order. These constraint equations provide general necessary conditions to make some approximation or simplification in some specific cases. When only three vector functions involved, we present a specific solution for the collision terms in the formalism of GCKT. We further specialize this solution in the form of relaxation-time approximation and present the self-consistent quantum kinetic theory at relaxation-time approximation which aligns with the self-consistent constraints. We find the quantum kinetic theory at relaxation-time approximation can be greatly simplified by defining or decomposing the distribution functions in the comoving fluid frame with velocity $u^{\mu}$. Some Wigner functions exhibit even no explicit dependence on the collision terms and the implicit dependence is totally coded into the distribution functions when we write the equations or functions in the fluid comoving frame.
@@
Although we didn’t take into account the mean electromagnetic field during deriving the quantum kinetic equation, Wigner functions, and the constraint conditions with collision terms, it is straightforward and easy to obtain the results when a mean field is involved. As we mentioned in the section II, the collision functions discussed in this work actually contain all possible terms, including mean field contribution. At mean field approximation, all the constraint conditions derived in previous sections will be satisfied automatically as verified in Gao:2018wmr ; Gao:2019znl ; Ma:2022ins . Therefore, we can simply decompose the general electromagnetic field in (5) into mean field part and quantized field part, separate the mean field part from the quantized field, and directly obtain the quantum kinetic equation and Wigner functions with mean electromagnetic field involved. The constraint conditions for collision terms from the quantized field part remain unchanged.
@@
We have presented the quantum kinetic equations and Wigner equations in 8-dimensional phase space, 4-dimensional coordinate space $x^{\mu}$ plus 4-dimensional momentum space $p^{\mu}$. In 8-dimensional phase space, the on-shell Dirac delta function is always involved. We can integrate the time-like component of momentum to eliminate the singular Dirac delta function and obtain the quantum kinetic theory in 7-dimensional phase space, which can be applied to make numerical calculation directly. Actually, it is a trivial task to obtain the 7-dimensional quantum kinetic theory from the 8-dimensional one in sections V and VI because only the onshell Dirac delta functions are involved and there are no derivative terms with respect to the momentum when the background electromagnetic field is absent.